OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope 9/7 and containing the midpoint of the segment whose end points are (2, -3) and (-6, 5). Place the symbols on the grid to complete the equation.

OpenStudy (anonymous):

you can get the midpoint [(2,-3)+(-6,5)]/2=(-2,1) so y-1=(9/7)(x+2)